OpenStudy (anonymous):

Clear system of fractions, then solve by substitution. 1/5x + 1/3y=1/15 & 1/5x +2y=8/5

hero (hero):

I assume it is this: \(\dfrac{1}{5}x + \dfrac{1}{3}y = \dfrac{1}{15}\) \(\dfrac{1}{5}x + 2y = \dfrac{8}{5}\)